# Modewise Additive Factor Model for Matrix Time Series

### v1

@@ 1 Introduction
Factor analysis is a fundamental tool for understanding common dependence among high-dimensional observations. Classical vector factor models provide a parsimonious representation of co-movements through a small number of latent factors, treating each observation as a long vector driven by shared components. These models have been rigorously developed for vector-valued time series in statistics, economics and finance, with seminal works (connor1986performance; bai2002determining; bai2003inferential; forni2004generalized; forni2005generalized; lam2011estimation; lam2012factor; fan2016projected) establishing consistency and inferential theory as both sample size and dimension grow. However, when observations possess natural matrix or tensor structure, such as panels of units and variables, spatiotemporal measurements, or import-export volumes across product categories, vector factor models fail to capture important multi-way dependencies (wang2019factor). Vectorizing a matrix or tensor severs inherent row and column relationships, discarding structural information and often yielding less efficient and less interpretable estimates.
This limitation has motivated substantial recent interest in matrix- and tensor-valued factor models. By preserving multi-way structure, these models capture within- and across-mode dependencies that would be lost through flattening, often yielding more interpretable factors and improved estimation efficiency. A prominent class of such models employs a multiplicative architecture, where a latent factor tensor interacts with loading matrices along each mode. For a matrix time series $\{\boldsymbol{X}_{t}\}_{t=1}^{n}$, a prototypical example is the Tucker factor model (wang2019factor; chen2020constrained; yu2022projected; yu2024dynamic):
@@
where $\boldsymbol{F}_{t}\in\mathbb{R}^{r_{1}\times r_{2}}$ is a matrix of latent factors and $\boldsymbol{A}\in\mathbb{R}^{d_{1}\times r_{1}},\boldsymbol{B}\in\mathbb{R}^{d_{2}\times r_{2}}$ are row and column loading matrices. Entry-wise, $X_{t,ij}=\boldsymbol{a}_{i}^{\top}\boldsymbol{F}_{t}\boldsymbol{b}_{j}+E_{t,ij}$, so the latent factors act globally across all rows and columns. This Tucker factor model framework extends naturally to tensor time series (chen2022factor; han2022rank; han2024tensor; babii2025tensor; chen2024semi; chen2024rank). A related but more constrained variant is the CP factor model (chang2023modelling; chang2024identification; chen2026estimation; han2024cp; chen2025diffusion; bolivar2025threshold), which assumes a superdiagonal latent factor tensor.
@@
To overcome this limitation, lam2025matrix introduced a main-effect Tucker factor model with spectral decomposition-based estimation, while yuan2023twoway proposed a two-way dynamic factor model using quasi-maximum likelihood estimation. In this paper, we introduce and analyze a Modewise Additive Factor Model (MAFM) for matrix time series that separately captures row-specific and column-specific latent effects. Each observation $\boldsymbol{X}_{t}\in\mathbb{R}^{d_{1}\times d_{2}}$ decomposes into a row-factor component, a column-factor component, and noise:
@@
In this paper, we develop a computationally efficient two-stage procedure for MAFM. The first stage, Modewise Inner-product Eigendecomposition (MINE), computes eigendecompositions of the modewise sample covariance matrices of $\boldsymbol{X}_{t}$ to obtain initial estimates. While the signal from one mode induces structural bias when estimating the other, this bias has weaker signal strength than the target component and can be controlled. The second stage, COMplement-Projected Alternating Subspace Estimation (COMPAS), iteratively refines these estimates through orthogonal complement projections, a key methodological innovation. Unlike the orthogonal projections in Tucker and CP factor models (han2024tensor; han2024cp), our orthogonal complement projections completely eliminate cross-modal interference: when estimating the loading matrix for one mode, projecting onto the orthogonal complement perfectly removes the signal from the other mode. Our theoretical analysis provides details of these improvements.
@@
Beyond estimation, we establish asymptotic normality of the estimated factor loading matrices and develop consistent estimators for variance-covariance components under proper conditions, yielding a data-driven statistical inference framework which is an aspect largely underexplored in the factor model literature. These results enable valid confidence interval construction and hypothesis testing, tools that are essential for empirical applications. Our inferential approach, based on spectral representations of the estimators, differs fundamentally from techniques used in vector factor models (bai2003inferential) and Tucker factor models (chen2023statistical; yu2022projected). Our approach can be extended naturally to other matrix and tensor factor model problems, such as inference of factor loading matrices using lagged autocovariance matrices (han2024tensor; chang2023modelling). Additionally, we derive sharp tail probability inequalities for quadratic forms of matrix time series under exponential-type tail conditions and temporal dependence, which may be of independent interest for the analysis of matrix and tensor time series models.
@@ 1.1 Notation
For a matrix $\boldsymbol{A}\in\mathbb{R}^{p_{1}\times p_{2}}$ of rank $r$, let its singular value decomposition (SVD) be $\boldsymbol{A}=\boldsymbol{U}_{1}\boldsymbol{\Lambda}\boldsymbol{U}_{2}^{\top}$, where $\boldsymbol{U}_{1}\in\mathbb{O}^{p_{1}\times r}$ and $\boldsymbol{U}_{2}\in\mathbb{O}^{p_{2}\times r}$ are the left and right singular matrices, and $\boldsymbol{\Lambda}={\rm diag}(\sigma_{1}(\boldsymbol{A}),\ldots,\sigma_{r}(\boldsymbol{A}))$ contains the singular values in descending order: $\sigma_{1}(\boldsymbol{A})\geq\cdots\geq\sigma_{r}(\boldsymbol{A})>0$. We write $\sigma_{\max}(\boldsymbol{A})=\sigma_{1}(\boldsymbol{A})$ and $\sigma_{\min}(\boldsymbol{A})=\sigma_{r}(\boldsymbol{A})$. The projection matrices onto the left and right singular subspaces are $\boldsymbol{P}_{\boldsymbol{A}}=\boldsymbol{U}_{1}\boldsymbol{U}_{1}^{\top}$ and $\boldsymbol{P}_{\boldsymbol{A}^{\top}}=\boldsymbol{U}_{2}\boldsymbol{U}_{2}^{\top}$. The spectral and Frobenius norms are denoted as $\|\boldsymbol{A}\|_{2}=\|\boldsymbol{A}\|=\sigma_{1}(\boldsymbol{A})$ and $\|\boldsymbol{A}\|_{\mathrm{F}}$, respectively. For two matrices $\boldsymbol{U},\widehat{}\boldsymbol{U}\in\mathbb{O}^{p\times r}$, let $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$ be the singular values of $\boldsymbol{U}^{\top}\widehat{}\boldsymbol{U}$. A natural measure of distance between their column spaces is $\|\widehat{}\boldsymbol{U}\widehat{}\boldsymbol{U}^{\top}-\boldsymbol{U}\boldsymbol{U}^{\top}\|_{2}=\sqrt{1-\sigma_{r}^{2}},$ which equals the sine of the largest principal angle between the column spaces.
@@ 2 Identifiability of Modewise Additive Factor Model
yuan2023twoway consider a similar factor model for high-dimensional matrix time series but impose substantially stronger assumptions for model identification. In addition to uncorrelatedness between row and column factors (C1), their framework requires: (i) mutual uncorrelatedness across rows of each factor matrix, i.e., $\text{Cov}(\mathbf{f}_{t,i},\mathbf{f}_{t,j})=0$ and $\text{Cov}(\boldsymbol{g}_{t,i},\boldsymbol{g}_{t,j})=0$ for all $i\neq j$; (ii) identical covariance and autocovariance structures across rows of each factor matrix, i.e., $\text{Cov}(\mathbf{f}_{t,i},\mathbf{f}_{t^{\prime},i})$ and $\text{Cov}(\boldsymbol{g}_{t,j},\boldsymbol{g}_{t^{\prime},j})$ are constant across $i,j$ for all $t,t^{\prime}\in[n]$; and (iii) diagonal covariance matrices $\text{Cov}(\mathbf{f}_{t,i})$ and $\text{Cov}(\boldsymbol{g}_{t,j})$. These assumptions are restrictive and atypical in both time series analysis and factor modeling. In contrast, our model identification relies on considerably milder conditions. As we demonstrate below, our estimation procedure also operates under standard regularity conditions that are broadly applicable.
@@ 3 Estimation
The orthogonal complement projection eliminates the row-factor component $\widetilde{\boldsymbol{F}}_{t}\boldsymbol{U}_{\boldsymbol{A}}^{\top}$, while preserving the column-factor component. Under proper conditions on the projected noise matrix $\boldsymbol{E}_{t}\boldsymbol{U}_{\boldsymbol{A}\perp}$, estimation of the loading matrix $\boldsymbol{U}_{\boldsymbol{B}}$ based on $\boldsymbol{X}_{t}\boldsymbol{U}_{\boldsymbol{A}\perp}$ can be made significantly more accurate. A symmetric argument applies to estimating $\boldsymbol{U}_{\boldsymbol{A}}$ via multiplying $\boldsymbol{X}_{t}^{\top}$ by $\boldsymbol{U}_{\boldsymbol{B}\perp}$. We note that since $r_{1}\ll d_{2}$ and $r_{2}\ll d_{1}$, unlike orthogonal projections in Tucker and CP factor models (han2024tensor; han2024cp), the orthogonal complement projection does not reduce the noise variation.
@@ 4.1 Convergence Rates for Factor Loading Estimation
Assumption 1 parallels noise conditions in lam2011estimation; lam2012factor; han2024tensor; han2024cp; chen2022factor. It accommodates general dependence patterns among individual time series while maintaining analytical tractability. The Gaussian assumption, imposed for technical convenience, ensures fast convergence rates in our analysis. In principle, this could be relaxed to sub-Gaussian or heavier-tailed distributions. However, such generalizations would substantially complicate the formulas, conditions, and statistical outcomes in our time series framework without yielding additional insights. To maintain focus on the core contributions, we adopt the Gaussian error assumption, which simplifies exposition without compromising the fundamental methodology or findings.
@@
Assumption 2 accommodates a broad class of time series models, including causal ARMA processes with continuously distributed innovations (tsay2010analysis; tsay2018nonlinear; fan2003nonlinear). For notational simplicity, we impose a common $\theta_{1}$ for both factor processes, though different values could be accommodated without altering the proofs. The independence between factor processes $\boldsymbol{F}_{t}$ and $\boldsymbol{G}_{t}$ is a structural assumption that ensures separation: row factors drive across-column variation, while column factors drive across-row variation, without overlap.
@@
Assumption 5 is similar to the signal strength condition in lam2012factor; han2024tensor; chen2026estimation, and aligns with the pervasive condition on factor loadings (stock2002forecasting; bai2003inferential). It plays a key role in identifying the common factors and idiosyncratic noises in (2). The indices $\delta_{0},\delta_{1}$ measure factor loading strength, or the rate of signal strength growth as the dimension $d_{k}$ increases: $\delta_{0}$ characterizes the strength of the strongest factors, while $\delta_{1}$ corresponds to the weakest. When $\delta_{0}=\delta_{1}=0$, the factors are called strong factors; otherwise, the factors are called weak factors. For notational simplicity, we adopt common values of $\delta_{0}$ and $\delta_{1}$ across both dimensions, $d_{1}$ and $d_{2}$, although different values can be accommodated without affecting the analysis. The gap $\delta_{1}-\delta_{0}$ further characterizes the heterogeneity in loading strengths.
@@
The error bounds in Theorem 4.1 comprise three distinct components. The first term corresponds to the average of the noise-by-noise cross-products. If all factors are strong factors with $\delta_{0}=\delta_{1}=0$, this term reduces to order $1/d_{2}$ or $1/d_{1}$, consistent with non-iterative Tucker factor models (chen2023statistical). This rate illustrates the “blessing of dimensionality” (barigozzi2023quasi): larger cross-sectional dimensions help average out noise. The second term arises from the average of the signal-by-noise cross-products. Under strong factors, the order of rates $(nd_{1})^{-1/2}$ and $(nd_{2})^{-1/2}$ also matches Tucker factor models (han2024tensor). The third term captures cross-modal interference, reflecting the interaction between row and column factors as well as contamination from the irrelevant factor mode.
@@
The key insight in analyzing COMPAS (Algorithm 2) is that each iteration achieves error contraction, with contraction rates $\rho_{1}$ and $\rho_{2}$ for estimating $\boldsymbol{U}_{\boldsymbol{A}}$ and $\boldsymbol{U}_{\boldsymbol{B}}$, respectively. Unlike Tucker and CP factor models (han2024tensor; han2024cp), COMPAS does not reduce noise variation; instead, it eliminates only the third term in the initialization bounds (6) and (7), which represents cross-modal interference. The contraction factors $\eta_{\boldsymbol{A}}$ and $\eta_{\boldsymbol{B}}$ govern the vanishing rate of this term, rather than affecting the oracle accuracy of the loading space estimators as in han2024tensor. This reflects the role of alternating orthogonal complement projection: each mode benefits from refined estimates of the other, so that the error $\|\widehat{\boldsymbol{U}}_{\boldsymbol{A}}\widehat{\boldsymbol{U}}_{\boldsymbol{A}}^{\top}-\boldsymbol{U}_{\boldsymbol{A}}\boldsymbol{U}_{\boldsymbol{A}}^{\top}\|$ includes the leading error term of $\widehat{\boldsymbol{U}}_{\boldsymbol{B}}$ scaled by $\eta_{\boldsymbol{A}}$, and vice versa.
@@
Consider $d_{1}\asymp d_{2}$ with fixed ranks $r_{1},r_{2}$. Using covariance tensors (rather than auto-covariance tensors), CP factor models (han2024cp; chen2026estimation) achieve loading vectors error bounds of order $O(d_{1}^{-2+2\delta_{1}}+n^{-1/2}d_{1}^{-1/2+2\delta_{1}})$, while Tucker factor models (han2024tensor) achieve loading space error bounds $O(d_{1}^{-2+2\delta_{1}}+n^{-1/2}d_{1}^{-1/2-\delta_{0}+2\delta_{1}})$. Compared with the rates of our MAFM $O(d_{1}^{-1+\delta_{1}}+n^{-1/2}d_{1}^{-1/2-\delta_{0}/2+\delta_{1}})$ from Theorem 4.2, the first noise-by-noise error term in our bounds is always worse than in CP and Tucker factor models. For the second signal-by-noise error term, CP factor models are always worse than ours, while the comparison with Tucker factor models depends on the sign of $\delta_{0}/2-\delta_{1}$. Thus, convergence rates alone do not distinguish among these model classes. Although our initialization resembles the TIPUP initialization in chen2022factor; han2024tensor, the distinct model structures and projection designs can be exploited for model selection.
yuan2023twoway propose a similar two-way dynamic factor model restricted to strong factors ($\delta_{0}=\delta_{1}=0$). Their error bounds yield convergence rate $O_{\mathbb{P}}(d_{1}^{-1}+d_{1}^{-1/2}n^{-1/2})$, matching ours under the strong factor setting. However, their results are primarily asymptotic, whereas our analysis employs weaker conditions and provides non-asymptotic guarantees that extend to weak factor settings.

### v2

@@ 1 Introduction
Factor analysis is a fundamental tool for understanding common dependence among high-dimensional observations. Classical vector factor models provide a parsimonious representation of co-movements through a small number of latent factors, treating each observation as a long vector driven by shared components. These models have been rigorously developed for vector-valued time series in statistics, economics and finance, with seminal works (Connor and Korajczyk, 1986; Bai and Ng, 2002; Bai, 2003; Forni et al., 2004, 2005; Lam et al., 2011; Lam and Yao, 2012; Fan et al., 2016) establishing consistency and inferential theory as both sample size and dimension grow. However, when observations possess natural matrix or tensor structure, such as panels of units and variables, spatiotemporal measurements, or import-export volumes across product categories, vector factor models fail to capture important multi-way dependencies (Wang et al., 2019; Chen and Fan, 2023). Vectorizing a matrix or tensor severs inherent row and column relationships, discarding structural information and often yielding less efficient and less interpretable estimates.
This limitation has motivated substantial recent interest in matrix- and tensor-valued factor models. By preserving multi-way structure, these models capture within- and across-mode dependencies that would be lost through flattening, often yielding more interpretable factors and improved estimation efficiency. A prominent class of such models employs a multiplicative architecture, where a latent factor tensor interacts with loading matrices along each mode. For a matrix time series $\{\boldsymbol{X}_{t}\}_{t=1}^{n}$, a prototypical example is the Tucker factor model (Wang et al., 2019):
@@
where $\boldsymbol{F}_{t}\in\mathbb{R}^{r_{1}\times r_{2}}$ is a matrix of latent factors and $\boldsymbol{A}\in\mathbb{R}^{d_{1}\times r_{1}},\boldsymbol{B}\in\mathbb{R}^{d_{2}\times r_{2}}$ are row and column loading matrices. Entry-wise, $X_{t,ij}=\boldsymbol{a}_{i}^{\top}\boldsymbol{F}_{t}\boldsymbol{b}_{j}+E_{t,ij}$, so the latent factors act globally across all rows and columns. Recent research has extended this model and its estimation to more refined methodologies, including iterative refinement, formal inference, and more realistic matrix time-series settings such as time-varying and threshold models (Chen et al., 2020; Chen and Fan, 2023; Yu et al., 2022, 2024; Chen et al., 2025b, 2024a; Liu and Chen, 2022). Further, this Tucker factor model framework extends naturally to tensor time series (Chen et al., 2022; Han et al., 2022, 2024a; Babii et al., 2025; Chen et al., 2024b; Chen and Lam, 2024), while a related but more constrained variant is the CP factor model (Chang et al., 2023, 2024; Chen et al., 2026; Han et al., 2024b; Chen et al., 2025a; Bolivar et al., 2025), which assumes a superdiagonal latent factor tensor.
@@
To overcome this limitation, Lam and Cen (2025) introduced a main-effect Tucker factor model with spectral decomposition-based estimation, while Yuan et al. (2023) proposed a two-way dynamic factor model using quasi-maximum likelihood estimation. In this paper, we introduce and analyze a Modewise Additive Factor Model (MAFM) for matrix time series that separately captures row-specific and column-specific latent effects. Each observation $\boldsymbol{X}_{t}\in\mathbb{R}^{d_{1}\times d_{2}}$ decomposes into a row-factor component, a column-factor component, and noise:
@@
In this paper, we develop a computationally efficient two-stage procedure for MAFM. The first stage, Modewise Inner-product Eigendecomposition (MINE), computes eigendecompositions of the modewise sample covariance matrices of $\boldsymbol{X}_{t}$ to obtain initial estimates. While the signal from one mode induces structural bias when estimating the other, this bias has weaker signal strength than the target component and can be controlled. The second stage, COMplement-Projected Alternating Subspace Estimation (COMPAS), iteratively refines these estimates through orthogonal complement projections, a key methodological innovation. Unlike the orthogonal projections in Tucker and CP factor models (Han et al., 2024a, b), our orthogonal complement projections completely eliminate cross-modal interference: when estimating the loading matrix for one mode, projecting onto the orthogonal complement perfectly removes the signal from the other mode. Our theoretical analysis provides details of these improvements.
@@
Beyond estimation, we establish asymptotic normality of the estimated factor loading matrices and develop consistent estimators for variance-covariance components under proper conditions, yielding a data-driven statistical inference framework which is an aspect largely underexplored in the factor model literature. These results enable valid confidence interval construction and hypothesis testing, tools that are essential for empirical applications. Our inferential approach, based on spectral representations of the estimators, differs fundamentally from techniques used in vector factor models (Bai, 2003) and Tucker factor models (Chen and Fan, 2023; Yu et al., 2022). Our approach can be extended naturally to other matrix and tensor factor model problems, such as inference of factor loading matrices using lagged autocovariance matrices (Han et al., 2024a; Chang et al., 2023). Additionally, we derive sharp tail probability inequalities for quadratic forms of matrix time series under exponential-type tail conditions and temporal dependence, which may be of independent interest for the analysis of matrix and tensor time series models.
@@ 1.1 Notation
For a matrix $\boldsymbol{A}\in\mathbb{R}^{p_{1}\times p_{2}}$ of rank $r$, let its singular value decomposition (SVD) be $\boldsymbol{A}=\boldsymbol{U}_{1}\boldsymbol{\Lambda}\boldsymbol{U}_{2}^{\top}$, where $\boldsymbol{U}_{1}\in\mathbb{O}^{p_{1}\times r}$ and $\boldsymbol{U}_{2}\in\mathbb{O}^{p_{2}\times r}$ are the left and right singular matrices, and $\boldsymbol{\Lambda}={\rm diag}(\sigma_{1}(\boldsymbol{A}),\ldots,\sigma_{r}(\boldsymbol{A}))$ contains the singular values in descending order: $\sigma_{1}(\boldsymbol{A})\geq\cdots\geq\sigma_{r}(\boldsymbol{A})>0$. We write $\sigma_{\max}(\boldsymbol{A})=\sigma_{1}(\boldsymbol{A})$ and $\sigma_{\min}(\boldsymbol{A})=\sigma_{r}(\boldsymbol{A})$. The projection matrices onto the left and right singular subspaces are $\boldsymbol{P}_{\boldsymbol{A}}=\boldsymbol{U}_{1}\boldsymbol{U}_{1}^{\top}$ and $\boldsymbol{P}_{\boldsymbol{A}^{\top}}=\boldsymbol{U}_{2}\boldsymbol{U}_{2}^{\top}$. The spectral and Frobenius norms are denoted as $\|\boldsymbol{A}\|_{2}=\|\boldsymbol{A}\|=\sigma_{1}(\boldsymbol{A})$ and $\|\boldsymbol{A}\|_{\mathrm{F}}$, respectively. For two orthonormal matrices $\boldsymbol{U},\widehat{}\boldsymbol{U}\in\mathbb{O}^{p\times r}$, let $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$ be the singular values of $\boldsymbol{U}^{\top}\widehat{}\boldsymbol{U}$. A natural measure of distance between their column spaces is $\|\widehat{}\boldsymbol{U}\widehat{}\boldsymbol{U}^{\top}-\boldsymbol{U}\boldsymbol{U}^{\top}\|_{2}=\sqrt{1-\sigma_{r}^{2}},$ which equals the sine of the largest principal angle between the column spaces.
@@ 2 Identifiability of Modewise Additive Factor Model
Yuan et al. (2023) consider a similar factor model for high-dimensional matrix time series but impose substantially stronger assumptions for model identification. In addition to uncorrelatedness between row and column factors (C1), their framework requires: (i) mutual uncorrelatedness across rows of each factor matrix, i.e., $\text{Cov}(\mathbf{f}_{t,i},\mathbf{f}_{t,j})=0$ and $\text{Cov}(\boldsymbol{g}_{t,i},\boldsymbol{g}_{t,j})=0$ for all $i\neq j$; (ii) identical covariance and autocovariance structures across rows of each factor matrix, i.e., $\text{Cov}(\mathbf{f}_{t,i},\mathbf{f}_{t^{\prime},i})$ and $\text{Cov}(\boldsymbol{g}_{t,j},\boldsymbol{g}_{t^{\prime},j})$ are constant across $i,j$ for all $t,t^{\prime}\in[n]$; and (iii) diagonal covariance matrices $\text{Cov}(\mathbf{f}_{t,i})$ and $\text{Cov}(\boldsymbol{g}_{t,j})$. These assumptions are restrictive and atypical in both time series analysis and factor modeling. In contrast, our model identification relies on considerably milder conditions. As we demonstrate below, our estimation procedure also operates under standard regularity conditions that are broadly applicable.
@@ 3 Estimation
The orthogonal complement projection eliminates the row-factor component $\widetilde{\boldsymbol{F}}_{t}\boldsymbol{U}_{\boldsymbol{A}}^{\top}$, while preserving the column-factor component. Under proper conditions on the projected noise matrix $\boldsymbol{E}_{t}\boldsymbol{U}_{\boldsymbol{A}\perp}$, estimation of the loading matrix $\boldsymbol{U}_{\boldsymbol{B}}$ based on $\boldsymbol{X}_{t}\boldsymbol{U}_{\boldsymbol{A}\perp}$ can be made significantly more accurate. A symmetric argument applies to estimating $\boldsymbol{U}_{\boldsymbol{A}}$ via multiplying $\boldsymbol{X}_{t}^{\top}$ by $\boldsymbol{U}_{\boldsymbol{B}\perp}$. We note that since $r_{1}\ll d_{2}$ and $r_{2}\ll d_{1}$, unlike orthogonal projections in Tucker and CP factor models (Han et al., 2024a, b), the orthogonal complement projection does not reduce the noise variation.
@@ 4.1 Convergence Rates for Factor Loading Estimation
Assumption 1 parallels noise conditions in Lam et al. (2011); Lam and Yao (2012); Han et al. (2024a, b); Chen et al. (2022). It accommodates general dependence patterns among individual time series while maintaining analytical tractability. The Gaussian assumption, imposed for technical convenience, ensures fast convergence rates in our analysis. In principle, this could be relaxed to sub-Gaussian or heavier-tailed distributions. However, such generalizations would substantially complicate the formulas, conditions, and statistical outcomes in our time series framework without yielding additional insights. To maintain focus on the core contributions, we adopt the Gaussian error assumption, which simplifies exposition without compromising the fundamental methodology or findings.
@@
Assumption 2 accommodates a broad class of time series models, including causal ARMA processes with continuously distributed innovations (Tsay, 2010; Tsay and Chen, 2018; Fan and Yao, 2003). For notational simplicity, we impose a common $\theta_{1}$ for both factor processes, though different values could be accommodated without altering the proofs. The independence between factor processes $\boldsymbol{F}_{t}$ and $\boldsymbol{G}_{t}$ is a structural assumption that ensures separation: row factors drive across-column variation, while column factors drive across-row variation, without overlap.
@@
Assumption 5 is similar to the signal strength condition in Lam and Yao (2012); Han et al. (2024a); Chen et al. (2026), and aligns with the pervasive condition on factor loadings (Stock and Watson, 2002; Bai, 2003). It plays a key role in identifying the common factors and idiosyncratic noises in (2). The indices $\delta_{0},\delta_{1}$ measure factor loading strength, or the rate of signal strength growth as the dimension $d_{k}$ increases: $\delta_{0}$ characterizes the strength of the strongest factors, while $\delta_{1}$ corresponds to the weakest. When $\delta_{0}=\delta_{1}=0$, the factors are called strong factors; otherwise, the factors are called weak factors. For notational simplicity, we adopt common values of $\delta_{0}$ and $\delta_{1}$ across both dimensions, $d_{1}$ and $d_{2}$, although different values can be accommodated without affecting the analysis. The gap $\delta_{1}-\delta_{0}$ further characterizes the heterogeneity in loading strengths.
@@
The error bounds in Theorem 4.1 comprise three distinct components. The first term corresponds to the average of the noise-by-noise cross-products. If all factors are strong factors with $\delta_{0}=\delta_{1}=0$, this term reduces to order $1/d_{2}$ or $1/d_{1}$, consistent with non-iterative Tucker factor models (Chen and Fan, 2023). This rate illustrates the “blessing of dimensionality” (Barigozzi, 2023): larger cross-sectional dimensions help average out noise. The second term arises from the average of the signal-by-noise cross-products. Under strong factors, the order of rates $(nd_{1})^{-1/2}$ and $(nd_{2})^{-1/2}$ also matches Tucker factor models (Han et al., 2024a). The third term captures cross-modal interference, reflecting the interaction between row and column factors as well as contamination from the irrelevant factor mode.
@@
The key insight in analyzing COMPAS (Algorithm 2) is that each iteration achieves error contraction, with contraction rates $\rho_{1}$ and $\rho_{2}$ for estimating $\boldsymbol{U}_{\boldsymbol{A}}$ and $\boldsymbol{U}_{\boldsymbol{B}}$, respectively. Unlike Tucker and CP factor models (Han et al., 2024a, b), COMPAS does not reduce noise variation; instead, it eliminates only the third term in the initialization bounds (6) and (7), which represents cross-modal interference. The contraction factors $\eta_{\boldsymbol{A}}$ and $\eta_{\boldsymbol{B}}$ govern the vanishing rate of this term, rather than affecting the oracle accuracy of the loading space estimators as in Han et al. (2024a). This reflects the role of alternating orthogonal complement projection: each mode benefits from refined estimates of the other, so that the error $\|\widehat{\boldsymbol{U}}_{\boldsymbol{A}}\widehat{\boldsymbol{U}}_{\boldsymbol{A}}^{\top}-\boldsymbol{U}_{\boldsymbol{A}}\boldsymbol{U}_{\boldsymbol{A}}^{\top}\|$ includes the leading error term of $\widehat{\boldsymbol{U}}_{\boldsymbol{B}}$ scaled by $\eta_{\boldsymbol{A}}$, and vice versa.
@@
Consider $d_{1}\asymp d_{2}$ with fixed ranks $r_{1},r_{2}$. Using covariance tensors (rather than auto-covariance tensors), CP factor models (Han et al., 2024b; Chen et al., 2026) achieve loading vectors error bounds of order $O(d_{1}^{-2+2\delta_{1}}+n^{-1/2}d_{1}^{-1/2+2\delta_{1}})$, while Tucker factor models (Han et al., 2024a) achieve loading space error bounds $O(d_{1}^{-2+2\delta_{1}}+n^{-1/2}d_{1}^{-1/2-\delta_{0}+2\delta_{1}})$. Compared with the rates of our MAFM $O(d_{1}^{-1+\delta_{1}}+n^{-1/2}d_{1}^{-1/2-\delta_{0}/2+\delta_{1}})$ from Theorem 4.2, the first noise-by-noise error term in our bounds is always worse than in CP and Tucker factor models. For the second signal-by-noise error term, CP factor models are always worse than ours, while the comparison with Tucker factor models depends on the sign of $\delta_{0}/2-\delta_{1}$. Thus, convergence rates alone do not distinguish among these model classes. Although our initialization resembles the TIPUP initialization in Chen et al. (2022); Han et al. (2024a), the distinct model structures and projection designs can be exploited for model selection.
Yuan et al. (2023) propose a similar two-way dynamic factor model restricted to strong factors ($\delta_{0}=\delta_{1}=0$). Their error bounds yield convergence rate $O_{\mathbb{P}}(d_{1}^{-1}+d_{1}^{-1/2}n^{-1/2})$, matching ours under the strong factor setting. However, their results are primarily asymptotic, whereas our analysis employs weaker conditions and provides non-asymptotic guarantees that extend to weak factor settings.